OpenStudy (anonymous):

Rex mixed x liters of fruit syrup with y liters of water to make a fruit punch. Three times the number of liters of water he mixed was 6 more than twice the number of liters of syrup he mixed. The total number of liters of syrup and water he mixed was 7 liters. Which graph best shows the number of liters of fruit syrup and water that Rex added to make the fruit punch?

OpenStudy (anonymous):

Rex mixed x liters of fruit syrup with y liters of water to make a fruit punch. Three times the number of liters of water he mixed was 6 more than twice the number of liters of syrup he mixed. The total number of liters of syrup and water he mixed was 7 liters. Which graph best shows the number of liters of fruit syrup and water that Rex added to make the fruit punch? graph of lines 3 y equals 2x plus 11 and y equals minus x plus 7. Lines intersect at ordered pair 2, 5 graph of lines 3 y equals 2 x minus 4 and y equals minus x plus 7. Lines intersect at ordered pair 5, 2 graph of lines 3 y equals 2 x plus 6 and y equals minus x plus 7. Lines intersect at ordered pair 3, 4 graph of lines 3y equals 2x plus 1 and y equals minus x plus 7. Lines intersect at ordered pair 4, 3

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@Destinyyyy

OpenStudy (anonymous):

@Elsa213

Elsa213 (elsa213):

You ask more questions then what you answer >.<

OpenStudy (anonymous):

true lol

Elsa213 (elsa213):

"x + y = 7 --> x = -y + 7 3y = 2x + 6 now sub -y + 7 in for x in the 2nd equation 3y = 2(-y + 7) + 6 3y = -2y + 14 + 6 3y = -2y + 20 3y + 2y = 20 5y = 20 y = 4 now sub 4 in for y in the 1st equation x + y = 7 x + 4 = 7 x = 7 - 4 x = 3 I do not see a graph, however, he added 3 liters of fruit syrup and 4 liters of water." http://inquus.org/updates/5261404ce4b0c23cd54a7ad9

OpenStudy (anonymous):

thank you so much

Elsa213 (elsa213):

np ;p

OpenStudy (anonymous):

your the best

Elsa213 (elsa213):

Ik i am

OpenStudy (anonymous):

lol;)